OpenStudy (anonymous):

maclaurin series cos(x^2) - 1

OpenStudy (sirm3d):

start with \[\cos y = \sum_{n=0}^\infty\frac{(-1)^n}{(2n)!}y^{2n}\] or \[\cos y = 1-\frac{y^2}{2!}+\frac{y^4}{4!}-\frac{y^6}{y!}+\cdots\]

OpenStudy (anonymous):

The McLaurin serie of the function cos x is : \[\Large \cos x= 1-\frac{x^2}{2!}+\frac{x^4}{4!}+\cdots+(-1)^p\frac{x^{2p}}{(2p)!}+\cdots\] So : \[\Large \cos x^2-1=-\frac{x^4}{2!}+\frac{x^8}{4!}+\cdots+(-1)^p\frac{x^{4p}}{(2p)!}+\cdots \]

OpenStudy (anonymous):

I need it in summation notation?

OpenStudy (anonymous):

What do I do with the remaining - 1?

OpenStudy (anonymous):

OK ! IT IS LIKE THIS : \[\Large \cos x^2-1=\sum_{k=1}^{+\infty}(-1)^k\frac{x^{4k}}{(2k)!}\]

OpenStudy (anonymous):

Does it matter is k starts at 0 or 1?

OpenStudy (anonymous):

It is matter, because we have "-1" so it begins from k=1

OpenStudy (anonymous):

Ok thank you so much! :)

OpenStudy (anonymous):

You are welcome !

OpenStudy (anonymous):

So it's for sure that and not